### v1

# Rethinking ASTE: A Minimalist Tagging Scheme Alongside Contrastive Learning

Abstract
Aspect Sentiment Triplet Extraction (ASTE) is a burgeoning subtask of fine-grained sentiment analysis, aiming to extract structured sentiment triplets from unstructured textual data. Existing approaches to ASTE often complicate the task with additional structures or external data. In this research, we propose a novel tagging scheme and employ a contrastive learning approach to mitigate these challenges. The proposed approach demonstrates comparable or superior performance in comparison to state-of-the-art techniques, while featuring a more compact design and reduced computational overhead. Notably, even in the era of Large Language Models (LLMs), our method exhibits superior efficacy compared to GPT 3.5 and GPT 4 in a few-shot learning scenarios. This study also provides valuable insights for the advancement of ASTE techniques within the paradigm of large language models.
Rethinking ASTE: A Minimalist Tagging Scheme Alongside Contrastive Learning
[…]
1 Introduction
Aspect Sentiment Triplet Extraction (ASTE) is an emerging fine-grained [^1] Generally, sentiment analysis can be based on three levels, namely, document-based, sentence-based, and aspect-based (Jing et al., 2021). sentiment analysis task Pontiki et al. (2014, 2015, 2016) aimed at identifying and extracting structured sentiment triplets Peng et al. (2020), defined as (Aspect, Opinion, Sentiment), from unstructured text. Specifically, an Aspect term refers to the subject of discussion, an Opinion term provides a qualitative assessment of the Aspect, and Sentiment denotes the overall sentiment polarity, typically taken from a three-level scale (Positive, Neutral, Negative). For instance, consider the sentence: “The battery life is good, but the camera is mediocre.” The extracted triplets are: (battery life, good, Positive) and (camera, mediocre, Neutral). Figure 1 illustrates another example. Recent methods Wu et al. (2020b); Jing et al. (2021); Chen et al. (2022); Zhang et al. (2022); Liang et al. (2023) commonly utilize Pretrained Language Models (PLMs) to encode input text. The powerful representational capacity of PLMs has greatly advanced the performance in this field, yet there is a tendency to employ complex classification head designs and leverage information enhancement techniques to achieve marginal performance improvements.
Figure 1: An example for the ASTE task illustrating Aspect terms in purple, Opinion terms with Negative sentiment in blue, and Opinion terms with Positive sentiment in orange.
In this work, we attribute the current challenges in ASTE to two main factors: 1) the longstanding overlook of the conical embedding distribution problem and 2) imprudent tagging scheme design. We critically examine the conventional 2D tagging method, commonly known as the table-filling approach, to reassess the efficacy of tagging schemes. Highlighting the critical role of tagging scheme optimization, we delve into what constitutes an ideal scheme for ASTE. We analyze the advantages of the full matrix approach over the half matrix approach and decompose the labels in the full matrix into 1) location and 2) classification. Thanks to this, we come up with a new tagging scheme with a minimum number of labels to effectively reduce the complexity of training and inference. Moreover, this tagging scheme can be well aligned with our novel contrastive learning mechanism. To the best of our knowledge, this is the first formal analysis of the tagging scheme to guide a more rational design and the first attempt to adopt token-level contrastive learning to improve the PLM representations’ distribution and facilitate the learning process.
The contributions of this work can be summarized as follows:
  • We offer the first critical evaluation of the 2D tagging scheme, particularly focusing on the table-filling method. This analysis pioneers in providing a structured framework for the rational design of tagging schemes.
  • We introduce a simplified tagging scheme with the least number of label categories to date, integrating a novel token-level contrastive learning approach to enhance PLM representation distribution.
  • Our study addresses ASTE challenges in the context of LLMs, developing a tailored in-context learning strategy. Through evaluations on GPT 3.5-Turbo and GPT 4, we establish our method’s superior efficiency and effectiveness.
2 Related Work
2.1 ASTE
Peng et al. (2020) proposes a pipeline method to bifurcate ASTE tasks into two stages, extracting (Aspect, opinion) pairs initially and predicting sentiment polarity subsequently. However, the error propagation hampers pipeline methods, rendering them vulnerable to end-to-end counterparts. End-to-end strategies, under the sequence-labeling approach, treat ASTE as a 1D “B-I-O” tagging scheme. ET Xu et al. (2020) introduces a position-aware tagging scheme with a conditional random field (CRF) module, effectively addressing span overlapping issues unhandled by Peng et al. (2020).
Recent advances in the end-to-end paradigm delicately grasp the peculiarity of ASTE tasks and come up with a proficient 2D table-filling tagging scheme. Wu et al. (2020a) pioneer the adoption of a grid tagging scheme (GTS) for ASTE, yielding substantial performance gains. While subsequent research refines and enhances GTS, it is not devoid of drawbacks. Instances involving multi-word Aspect/Opinion constructs risk relation inconsistency and boundary insensitivity Zhang et al. (2022). To overcome these, BDTF Zhang et al. (2022) designs a boundary-driven tagging scheme, effectively reducing boundary prediction errors. Alternative research augments GTS by integrating external semantic information as structured knowledge into their models. $\mathrm{S}^{3}\mathrm{E}^{2}$ Chen et al. (2021b) retains the GTS tagging scheme while introducing novel semantic and syntactic enhancement modules between word embedding outputs and the tagging scheme. EMGCN Chen et al. (2022) offers a distinct perspective, incorporating external knowledge from four aspects, namely, Part-of-Speech Combination, Syntactic Dependency Type, Tree-based Distance, and Relative Position Distance through an exogenous hard-encoding strategy. SyMux Fei et al. (2022) contributes a unified tagging scheme capable of all ABSA subtasks synthesizing insights from incorporating GCN, syntax encoder, and representation multiplexing.
2.2 Contrastive Learning
The idea of contrastive learning can be traced back to the 1990s Le-Khac et al. (2020) and was first formally proposed by Hadsell et al. (2006). While a unified definition of contrastive learning is still absent Le-Khac et al. (2020), its essence revolves around establishing proximity among similar instances and separation between dissimilar ones within a designated metric space Jaiswal et al. (2021). The pioneer work of Wang and Isola (2020) introduced the concepts of alignment and uniformity. Alignment quantifies the closeness between akin samples’ representations, while uniformity gauges the degree that representations are distributed evenly on a unit hypersphere. Note that a representation distribution satisfying alignment and uniformity is linearly separableWu et al. (2023) and facilitates the classification. Thereby, contrastive learning boosts represent learning by improving the alignment and uniformity of the representations. However, recent investigations indicate that representation distributions in pretrained models often diverge from these expectations. Liang et al. (2021) computed similarities for randomly sampled word pairs, revealing that word embeddings in an Euclidean space cluster within a confined cone, rather than uniformly distributed.
While contrastive learning has gained popularity in diverse NLP domains Wu et al. (2020b); Giorgi et al. (2021); Gao et al. (2021); Zhang et al. (2021), its application to ASTE remains relatively unexplored. Ye et al. (2021) adopts contrastive learning into triplet extraction in a generative fashion. Wang et al. (2022) takes contrastive learning as a data augmentation approach. Yang et al. (2023) proposed an enhancement approach in pairing with two separate encoders.
3 Method
Figure 2 presents our framework. In the training phase, the process unfolds as follows: first, the input sentence is encoded by a PLM encoder, such as BERT (Devlin et al., 2018) or RoBERTa (Liu et al., 2022). Next, a single-layer classification head predicts the tagging scheme using these embeddings, and simultaneously, a contrastive loss is calculated. Then, a focal loss is computed between the predicted and ground truth tagging schemes. Finally, the focal loss and contrastive loss are weighted into the total loss. By backpropagation and convergence of the total loss, it can be expected that the predicted tagging schemes will be more and more close to their ground truth.
Algorithm 1 provides a pseudo-code description for the training process of our proposed framework.
Figure 2: Schematic diagram of our proposed framework which contains: 1) a contrastive learning mechanism that aligns Aspect and Opinion; and 2) a tagging scheme encompassing 5 label classes: NULL, CTD, POS, NEU, and NEG.
3.1 Task Formulation
Given a sentence containing $|\mathcal{S}|$ tokens $\mathcal{S}_{|\mathcal{S}|}=(w_{1},w_{2},...,w_{|\mathcal{S}|})$, an ASTE algorithm should return all the triplets $\mathcal{T}_{|\mathcal{T}|}$, where $\mathcal{T}_{k}=(A_{k},O_{k},S_{k})$, $k\in\{1,2,...,|\mathcal{T}|\}$. $A_{k}$, $O_{k}$, $S_{k}$, and $|\mathcal{T}|$ denote the Aspect term, Opinion term, Sentiment polarity, and number of triplets contained in $\mathcal{S}$, respectively.
3.2 Pretrained Encoder
Given the input sentence $\mathcal{S}$ and pretrained model Encoder , it outputs the hidden word embeddings $\mathcal{H}_{|\mathcal{H}|}$:
$\mathcal{H}_{|\mathcal{H}|}=\texttt{Encoder}(\mathcal{S}_{|\mathcal{S}|}),$ (1)
where $\mathcal{H}_{|\mathcal{H}|}=(h_{1},h_{2},...,h_{|\mathcal{H}|})$.
3.3 Tagging Scheme
Rethinking the 2D tagging scheme:
Lemma 1. Specific to the ASTE task, when we take it as a 2D-labeling problem, we are to 1) find a set of tagging strategies to establish a 1-1 map between each triplet and its corresponding tagging matrix.
Lemma 2. In a 2D-tagging problem, there are at least three basic goals that must be met: 1) correctly identifying the (Aspect, Opinion) pairs, 2) correctly classifying the sentiment polarity of the pair based on the context, and 3) avoiding boundary errors, such as overlapping [^2] It occurs when one single word belongs to multiple classes in different triplets. , confusion [^3] It occurs when there is a lack of location restrictions so that multiple neighbored candidates can not be uniquely distinguished. , as per Lemma 1 and conflict [^4] It occurs when one single word is composed of multiple tokens, and the predict gives predictions that are not aligned with the word span. .
Property 1. From the above assumptions, it can be concluded that using fewer yet enough (that is, following the 1-1 map properties in Lemma 1, as well as avoiding the issues in Lemma 2) labels will make it a theoretically or heuristically better tagging scheme.
In essence, a 2D tagging scheme in ASTE serves two primary functions: locating and classifying. To develop an effective tagging scheme, three pitfalls must be avoided: overlapping, confusion, and conflict. Furthermore, building on this foundation, it is crucial to minimize the number of labels introduced to simplify the learning process.
With the above knowledge, our tagging scheme employs the full matrix so that the row-column relations can be seen as a directional graph, which naturally contains the order information. Taking advantage of this, we can see it as a coordinate system, where the row index indicates the Aspect term and the columns index indicates the Opinion. Then we can represent any triplet by a rectangular shadowed area in the matrix, which meets Lemma 3.3. Hereafter, this labeling can be taken as a “place holder”, which has supported a multi-to-one map, whilst weaker than the requirement in Lemma 3.3.
To satisfy all the conditions, we further introduce another tagging, “sentiment & beginning tag”. This tagging scheme specializes in recognizing the top-left corner of a “shadowed” area. Meanwhile, it takes a value from the sentiment polarity, i.e.Positive, Neutral, Negative. This tagging is crucial to both decide the boundary of an area and classify the sentiment polarity. It can be strictly proved that our tagging scheme perfectly meets the requirements and takes the least possible classes of labels.
Figure 3 shows a comprehensive case of our tagging scheme, in which the left matrix is an appearance of our tagging scheme, and it can be decomposed into two separate components. The middle matrix is the first component, which takes only one tag to locate the up-left beginning of an area, and the second component simply predicts a binary classification to figure out the full area.
When we add the two components together, we have the left tagging scheme, where the “Sentiment & Beginning Tag” is like a trigger (just like you click your mouse), and the “Place Holder” is like a “continued shift” (continue to hold and drag the mouse to the downright).
Note that, this design benefits the tagging scheme’s decode process. By scanning across the matrix, we only start an examination function when triggered by a beginning label like this, and then search by row and column until it meets any label except a “continued” (“CTD”).
Figure 3: Decomposition of the tagging scheme into two components: 1) a beginning mark matrix with sentiment labels; and 2) a placeholder matrix denoting regions of triplets with “1”s and default regions with “0”s. Remember that each row is taken as candidates for an Aspect and each column is taken as candidates for an Opinion. Naturally, each cell in the square matrix can be seen as an ordered pair for a unique candidate of <Aspect, Opinion>.
3.4 Token-level Contrastive Learning
The motivation to adopt contrastive learning is to improve the distribution of representations output by PLMs. The core idea lies in that, after fine-tuning on a specific task, the PLM encoder should embed words with similar roles to distribute closer and drive the different ones to be farther.
Specifically in the ASTE task, for example, we take each word into either Aspect, Opinion, or None of which. So, there should be at least 3 classes for sequence tagging. Our insight is to award the model to learn a sophisticated map to embed words to cluster with those sharing the same class, given a contextual input.
Inspired by (Schroff et al., 2015), we 1) take the Euclidean distance as a negative metric on the similarity and 2) introduce a margin $\mathcal{\alpha}$ to enforce the gap between two similar hidden word representations, that is, stop pulling $\mathcal{H}_{i}$ and $\mathcal{H}_{j}$ closer when there is $\left\|\mathcal{H}_{i}-\mathcal{H}_{j}\right\|^{2}\leq\alpha$ (avoiding similar representations from squeezing too much with each other).
So, the metric of similarity is
$\mathrm{Sim}_{i,j}=\mathbf{Sim}(\mathcal{H}_{i},\mathcal{H}_{j})=-\left\|\mathcal{H}_{i}-\mathcal{H}_{j}\right\|^{2}$ (2)
Hence, the similarities $\mathrm{Sim}_{i,j},i,j\in\{1,2,...,|\mathcal{H}|\}$ forms a similarity matrix $\mathbf{Sim}_{|H|\times|H|}$.
To make the representations closer among tokens within the same class and farther between that of different classes, while keeping a margin of $d$, we can simply maximize $\mathrm{Sim}_{i,j}$ if $\mathcal{H}_{i}$ and $\mathcal{H}_{j}$ shares the same class and else minimize $\mathrm{Sim}_{i,j}$.
Defining a matrix $\mathbf{M_{|\mathcal{H}|\times|\mathcal{H}|}}$, it controls whether to pull (closer) or push (farther) between the hidden representation of words, that is, the “Contrastive Mask” used to calculate the contrastive loss. In Figure 2, it is the left-side strict upper-triangle matrix in blue, orange and grey555 The meaning of colors: The blue cell means that on this cell the row representation is in a different class from the column representation, and the orange cell means that with same class. .
$\mathcal{L}_{contrastive}$ (3)
$=$ $\sum_{(\mathcal{H}_{i},\mathcal{H}_{j})^{+}}\max\{\left\|\mathcal{H}_{i}-\mathcal{H}_{j}\right\|^{2},\ d\}$
-- $\sum_{(\mathcal{H}_{i},\mathcal{H}_{j})^{-}}\min\{\left\|\mathcal{H}_{i}-\mathcal{H}_{j}\right\|^{2},\ d\}$
$=$ $\sum_{i=1}^{|\mathcal{H}|}\sum_{j=i+1}^{|\mathcal{H}|}\max\{\mathbf{Sim}(\mathcal{H}_{i},\mathcal{H}_{j}),\ d\}\cdot\mathbf{M}_{i,j}.$
$=$ $\sum_{i=1}^{|\mathcal{H}|}\sum_{j=1}^{|\mathcal{H}|}\big{(}\max\{\mathbf{Sim_{|\mathcal{H}|\times|\mathcal{H}|}},\ d\}\circ\mathbf{M_{|\mathcal{H}|\times|\mathcal{H}|}}\big{)}_{i,j}$
where $\circ$ is a notation for the Hadamard product Horn (1990).
Figure 4: A more detailed illustration of our contrastive learning mechanism.
Figure 4 depicts the workflow of our contrastive learning strategy and showcases the detail of the Contrastive Mask. Note that for the purpose of contrast learning, the order of each pair of words is redundant, so using the strict upper matrix is enough and we mask the lower triangle.
Noteworthily, a PLM encoder, like BERT, encodes the same word input to different vector representations when the context is different. For example, suppose two sentences, 1) “I really just like the old school plug-in ones”, and 2) “The old school is beautiful”. Explicitly the “school” should be labelled as Aspect in the former case whilst as Opinion in the latter. For these cases, a well-learned PLM encoder will encode the unique word “school” into different representations with accordance to the context.Therefore, given that a single word can have varying meanings and associated class labels depending on its context, its representations are inherently context-specific, differing accordingly. This characteristic allows for the direct application of contrastive learning to the hidden word representations without concern for contradiction, ensuring that contextual information is preserved without loss.
3.5 classification head
We employ a streamlined single-layer linear classification head, composed sequentially of a linear classification head, dropout, layer normalization, and a GELU activation function Hendrycks and Gimpel (2016), as follows:
$H_{1}=\texttt{LinearClsHead}(\mathcal{H}_{contrasted})$ (4)
$H_{3}=\texttt{LayerNorm}(H_{2})$
$\mathbf{Tag}_{pred}=\texttt{GELU}(H_{3})$
3.6 Focal Loss
The tagging matrix intrinsically contains more NULL labels than valid labels (POS, NEU, NEG, and CTD). This imbalance is exacerbated by the sentiment distribution, as demonstrated in Table 5 in the Appendix. Given its naturally sparse nature, this matrix faces challenges related to both sample imbalance and hard examples. To mitigate the pronounced imbalance between positive and negative samples in the training set, we employ the Focal LossLin et al. (2017), which approach prevents an inundation of easy negatives from overwhelming the training process. Our focal loss is defined as:
$\mathcal{L}_{focal}=\mathrm{FL}(p_{t})=-\alpha_{t}(1-p_{t})^{\gamma}\log(p_{t}),$ (5)
where $p_{t}$ is a brief notation for
$p_{t}=\begin{cases}p,&\textrm{if}\ y=1\\ 1-p,&\textrm{otherwise},\end{cases}$
$\alpha_{t}$ is a factor to balance between the positive and negative samples, $y$ denotes the ground-truth class and $p\in[0,1]$ represents the estimated probability for the class. $-\big{(}1-p_{t})^{\gamma}$ modulates the canonical cross entropy (CE) loss. Notably, when $\gamma=0$ and $\alpha_{t}=0.5$, the focal loss is equivalent to a standard cross entropy loss. The final loss function is a weighted sum of the contrastive loss and the focal loss:
$\mathcal{L}=\mathcal{L}_{focal}+\alpha\ \mathcal{L}_{contrastive},$ (6)
where $\alpha$ is the weighting hyperparameter.
Algorithm 1 Workflow of our framework.
Modules:
Input:
Raw sentences: $\mathcal{S}_{|\mathcal{S}|}$;
Ground truth triplets: $\mathcal{T}_{|\mathcal{T}|}^{gt}$ , where
$\mathcal{T}_{k}=(A_{k},O_{k},S_{k})$, $k\in\{1,2,...,|\mathcal{T}|\}$;
classes of contrasted labels: $\mathcal{C}$.
Output:
Predicted Triplets: $\mathcal{T}_{|\mathcal{T}|}^{pred}$;
[…]
Algorithm:
Repeat for $N$ epochs:
1:  Hidden word representation: $\mathcal{H}_{|\mathcal{H}|}=\texttt{PLMsEncoder}(\mathcal{S}_{|\mathcal{S}|})$;
2:  Tensor Operations: ||×$\mathcal{H}_{|\mathcal{H}|\times|\mathcal{H}|}=\texttt{expand}(\mathcal{H}_{|\mathcal{H}|})$, ||×$\mathcal{H}_{|\mathcal{H}|\times|\mathcal{H}|}^{T}=\mathcal{H}_{|\mathcal{H}|\times|\mathcal{H}|}.\texttt{transpose}()$;
3:  Similarity matrix: $\mathbf{Sim}_{|\mathcal{H}|\times|\mathcal{H}|}=-(\mathcal{H}_{|\mathcal{H}|\times|\mathcal{H}|}-\mathcal{H}_{|\mathcal{H}|\times|\mathcal{H}|}^{T})\circ(\mathcal{H}_{|\mathcal{H}|\times|\mathcal{H}|}-\mathcal{H}_{|\mathcal{H}|\times|\mathcal{H}|}^{T})$, where   $\mathbf{Sim}_{i,j}=-\left\|\mathcal{H}_{i}-\mathcal{H}_{j}\right\|^{2}$, and $\circ$ denotes the Hadamard product;
4:  Contrastive Mask matrix: $\mathbf{M}_{|\mathcal{H}|\times|\mathcal{H}|}$, where   $\mathbf{M}_{i,j}=1$ if $\mathcal{H}_{i},mathcal{H}_{j}\in\mathcal{C}_{p},p\in{1,2,3}$ else $-1$;
5:  Contrastive loss: $\mathcal{L}_{contrastive}=\newline \sum_{i=1}^{|\mathcal{H}|}\sum_{j=1}^{|\mathcal{H}|}\big{(}\mathbf{Sim_{|\mathcal{H}|\times|\mathcal{H}|}}\circ\mathbf{M_{|\mathcal{H}|\times|\mathcal{H}|}}\big{)}_{i,j};$
6:  Predicted tagging matrix: $\mathbf{Tag}_{|\mathcal{H}|\times|\mathcal{H}|}^{pred}=\texttt{ClsHead}(\mathcal{H}_{|\mathcal{H}|\times|\mathcal{H}|},\mathcal{H}_{|\mathcal{H}|\times|\mathcal{H}|}^{T})$;
7:  Focal loss: $\mathcal{L}_{focal}=\texttt{FocalLoss}(\mathbf{Tag}_{|\mathcal{H}|\times|\mathcal{H}|}^{pred},\mathbf{Tag}_{|\mathcal{H}|\times|\mathcal{H}|}^{gt})$;
8:  Weighted Loss: $\mathcal{L}=\mathcal{L}_{focal}+\alpha\mathcal{L}_{contrastive}$.
9:  Backward propagation.
Predicted triplets:
$\mathcal{T}_{|\mathcal{T}|}^{pred}=\texttt{TaggingDecoder}(\mathbf{Tag}_{|\mathcal{H}|\times|\mathcal{H}|}^{pred})$
Metric:
[…]
4 Experiments
4.1 Implementation Details
All experiments were conducted on a single RTX 2080 Ti. The best model weight on the development set is saved and then evaluated on the test set. For the PLM encoder, the pretrained weights bert_base_uncased and roberta_base are downloaded from (Wolf et al., 2020). GPT 3.5-Turbo and GPT 4 are implemented using OpenAI API . The learning rate is $1\times 10^{-5}$ for the PLM encoder, and $1\times 10^{-3}$ for the classification head.
4.2 Datasets
We evaluate our method on two canonical ASTE datasets derived from the SemEval Challenges Pontiki et al. (2014, 2015, 2016). These datasets serve as benchmarks in the majority of Aspect-based Sentiment Analysis (ABSA) research. The first dataset, denoted as $\mathcal{D}_{1}$, is the Aspect-oriented Fine-grained Opinion Extraction (AFOE) dataset introduced by Wu et al. (2020a). The second dataset, denoted as $\mathcal{D}_{2}$, is a refined version by Xu et al. (2020), building upon the work of Peng et al. (2020). Further details are provided in Table 5.
4.3 Baselines
We evaluate our method against various techniques including pipeline, sequence-labeling, MRC-based, table-filling and LLM-based approaches. Detailed descriptions of each method can be found in the Appendix Table 8.
Methods 14Res 14Lap 15Res 16Res
P R F1 P R F1 P R F1 P R F1
Pipeline
[…]
Sequence-tagging
Span-BART Yan et al. (2021) 65.52 64.99 65.25 61.41 56.19 58.69 59.14 59.38 59.26 66.60 68.68 67.62
JET Xu et al. (2020) 70.56 55.94 62.40 55.39 47.33 51.04 64.45 51.96 57.53 70.42 58.37 63.83
MRC-based
Dual-MRC Mao et al. (2021) 71.55 69.14 70.32 57.39 53.88 55.58 63.78 51.87 57.21 68.60 66.24 67.40
$\textrm{BMRC}^{\dagger}$ Chen et al. (2021a) 72.17 65.43 68.64 65.91 52.15 58.18 62.48 55.55 58.79 69.87 65.68 67.35
COM-MRC Zhai et al. (2022) 75.46 68.91 72.01 62.35 58.16 60.17 68.35 61.24 64.53 71.55 71.59 71.57
Triple-MRC Zou et al. (2024) - - 72.45 - - 60.72 - - 62.86 - - 68.65
Table-filling
GTS Wu et al. (2020a) 67.76 67.29 67.50 57.82 51.32 54.36 62.59 57.94 60.15 66.08 66.91 67.93
Double-encoder Jing et al. (2021) 67.95 71.23 69.55 62.12 56.38 59.11 58.55 60.00 59.27 70.65 70.23 70.44
EMC-GCN Chen et al. (2022) 71.21 72.39 71.78 61.70 56.26 58.81 61.54 62.47 61.93 65.62 71.30 68.33
BDTF Zhang et al. (2022) 75.53 73.24 74.35 68.94 55.97 61.74 68.76 63.71 66.12 71.44 73.13 72.27
STAGE-1D Liang et al. (2023) 79.54 68.47 73.58 71.48 53.97 61.49 72.05 58.23 64.37 78.38 69.10 73.45
STAGE-2D Liang et al. (2023) 78.51 69.3 73.61 70.56 55.16 61.88 72.33 58.93 64.94 77.67 68.44 72.75
STAGE-3D Liang et al. (2023) 78.58 69.58 73.76 71.98 53.86 61.58 73.63 57.9 64.79 76.67 70.12 73.24
DGCNAP Li et al. (2023) 72.90 68.69 70.72 62.02 53.79 57.57 62.23 60.21 61.19 69.75 69.44 69.58
LLM-based
GPT 3.5 zero-shot 44.88 55.13 49.48 30.04 41.04 34.69 36.02 53.40 43.02 39.92 57.78 47.22
GPT 3.5 few-shots 52.36 54.63 53.47 29.91 36.04 32.69 45.48 61.44 52.01 49.50 67.12 56.98
GPT 4 zero-shot 32.99 38.13 35.37 17.81 22.55 19.90 27.85 37.73 32.05 32.17 43.00 36.80
GPT 4 few-shots 47.25 49.20 48.20 26.04 33.64 29.35 39.94 51.13 44.85 43.72 54.86 48.66
Ours
ContrASTE 76.1 75.08 75.59 66.82 60.68 63.61 66.50 63.86 65.15 75.52 74.14 74.83
Table 1: Experimental results on $\mathcal{D}_{2}$ Xu et al. (2020). The best results are highlighted in bold, while the second best results are underlined.
4.4 Performance on ASTE Tasks
We evaluate ASTE performance using the widely accepted (Precision, Recall, F1) metrics. The result on dataset $\mathcal{D}_{2}$ can be found in Table 1 and on $\mathcal{D}_{1}$ is presented in Appendix Table 7. The best results are indicated in bold, while the second best results are underlined. Our proposed method consistently achieves state-of-the-art performance or ranks second across all evaluated cases.
Significantly, on dataset $\mathcal{D}_{1}$, proposed method achieves a substantial improvement of 3.08 percentage points in F1 score on the 14Lap subset. This improvement is particularly noteworthy considering that the best score on this dataset is the lowest among all the datasets, showcasing our ability to effectively handle challenging instances. Moreover, on the 14Res subset, our F1 score surpasses 76.00+, which, to the best of our knowledge, is the highest reported performance.
Turning to dataset $\mathcal{D}_{2}$, our method outperforms all state-of-the-art approaches by more than 1 percentage point on the 14Res, 14Lap, and 16Res subsets. Only on the 16Res subset, the BDTF method (Zhang et al., 2022) achieves a slightly better performance.
It is worth highlighting that, compared to previous methods, our approach achieves significantly higher recall in the majority of cases in both $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$, without significantly compromising precision. This observation underscores the robustness of our method in minimizing the omission of true triplets. This behavior can be attributed to our contrastive learning process, where the learned representations contribute to accurate predictions, thereby enhancing the effectiveness of triplet identification.
4.5 Performance on Other ABSA Tasks
Our method can also effectively handle other ABSA subtasks, including Aspect Extraction (AE), Opinion Extraction (OE), and Aspect Opinion Pair Extraction (AOPE). AE aims to extract all the (Aspect) terms, OE aims to extract all the Opinion terms, and AOPE aims to extract all the (Aspect, Opinion) pairs from raw text. The results for these tasks are presented in Table 6 in the Appendix, where our method consistently achieves best F1-scores across nearly all tasks.
Models $\mathcal{D}_{1}$ $\mathcal{D}_{2}$
14Res 14Lap 15Res 16Res 14Res 14Lap 15Res 16Res
ContrASTE 76.00 64.07 65.43 71.80 75.59 63.61 65.15 74.83
w/o. RoBERTa 74.12 63.18 62.95 69.41 72.66 62.15 63.25 70.71
    $\Delta$F_1 -1.88 -0.89 -2.48 -2.39 -2.93 -1.46 -1.90 -4.12
    w/o. contr 72.61 61.94 58.14 68.16 71.72 61.49 58.11 68.03
    $\Delta$F_1 -3.39 -2.13 -7.29 -3.64 -3.87 -2.12 -7.04 -6.80
    w/o. tag 67.78 54.98 60.75 62.62 65.83 54.98 58.73 67.63
    $\Delta$F_1 -8.22 -9.09 -4.68 -9.18 -9.76 -8.63 -6.42 -7.20
Table 2: Ablation study on F1, where “w/o. RoBERTa” denotes “Replace RoBERTa with bert-base-uncased”, “w/o. contr” denotes without the contrastive learning mechanism, and “w/o. tag” denotes “replace our tagging scheme with a baseline”.
5 Analysis
5.1 Ablation Study
Encoder. In the ablation experiment part, by replacing RoBERTa by BERT, The results obtained declined slightly while still yield most other methods.
Contrastive Learning. First, we deactivating the contrastive mechanism in our method (denoting “w/o. contr”) by setting the coefficient of the contrastive loss to 0. The results in Table 2 illustrate a significant F1-score decrease of $0.22\sim 2.93$ percentage points.
Tagging Scheme. Third, we substitute our proposed scheme with the conventional GTS tagging scheme Wu et al. (2020a), which results in a significant performance decline (Table 2) by $1.46\sim 6.14$ percentage points. This indicates that the contrastive learning methods, within our framework, is of strong reliance on an appropriate tagging scheme. This reinforces the effectiveness of our straightforward yet impactful tagging scheme.
5.2 Effect of Contrastive Learning
In Appendix Figure 5 an example is shown of how contrastive learning has improved the representation, where the left subplot is the tokens output by RoBERTa without contrastive learning, and the right one is with contrastive learning. Note that the Principal Component Analysis (PCA) (Maćkiewicz and Ratajczak, 1993) is adopted to reduce the dimensions of the vectors to 2.
5.3 Efficiency Analysis
Table 3: Efficiency Analysis
Model Memory Num Params Epoch Time Inf Time F1(%) Device
Span-ASTE 3.173GB - 108s - 71.62 Tesla v100
BDTF 8.103GB - 135s - 74.73 Tesla v100
GPT 3.5-Turbo $>80GB$ 175B - 0.83s 49.48 OpenAI API
GPT 4 $>80GB$ 1760B - 1.56s 35.37 OpenAI API
Ours 7.11GB 0.12B 10s 0.01s 76.00 2080 Ti
Table 4: Tagging Scheme Comparisons
Method Num Tags Linguistic Features Half/Full Matrix
GTS 6 None Half
Double-encoder 9 None Half
EMC-GCN 10 4 Groups Full
BDTF $2\times 2\times 3$ None Half
STAGE $2\times 2\times 4$ None Half
DGCNAP 6 POS-tagging Half
Ours $5$ None Full
Table 3 shows an efficiency analysis, where our method is in significantly higher efficiency. While our approach demonstrates efficient management of memory and parameter utilization, surpassing other methods in terms of runtime performance, it is crucial to note the unsatisfactory performance of LLM-based methods, as demonstrated by Table 7, Table 1 and Table 9. This observation highlights that employing a general LLM with a large number of parameters does not yield desirable results in specific tasks like ASTE, even with the assistance of few-shot memorizing processes. Moreover, the incorporation of LLM introduces a significant computational resource overhead in general application scenarios. Although fine-tuning the parameters of LLM itself may offer some improvement, there exists a risk of collapsive forgetting. Furthermore, Table 4 provides a comparative analysis of tagging schemes. Our method achieves its objective without the need for additional linguistic information and utilizes fewer taggings.
6 Conclusion
In this work, we have introduced an elegant and efficient framework for ASTE, achieving SOTA performance. Our approach is built upon two effective components: a new tagging scheme and a novel token-level contrastive learning implementation. The ablation study demonstrates the synergy between these components, reducing the need for complex model designs and external information enhancements. The limitation of this work lies in that there lefts potential to further optimize the framework by investigating various classification head strategies.
Appendix A Appendix
Datasets #S #A #O #S1 #S2 #S3 #T
14Res $\mathcal{D}_{1}$ Train 1259 1008 849 1456 164 446 2066
Dev 315 358 321 352 44 93 489
Test 493 591 433 651 59 141 851
$\mathcal{D}_{2}$ Train 1266 986 844 1692 166 480 2338
Dev 310 396 307 404 54 119 577
Test 492 579 437 773 66 155 994
14Lap $\mathcal{D}_{1}$ Train 899 731 693 691 107 466 1264
Dev 225 303 237 173 42 118 333
Test 332 411 330 305 62 101 468
$\mathcal{D}_{2}$ Train 906 733 695 817 126 517 1460
Dev 219 268 237 169 36 141 346
Test 328 400 329 364 63 116 543
15Res $\mathcal{D}_{1}$ Train 603 585 485 668 24 179 871
Dev 151 182 161 156 8 41 205
Test 325 353 307 293 19 124 436
$\mathcal{D}_{2}$ Train 605 582 462 783 25 205 1013
Dev 148 191 183 185 11 53 249
Test 322 347 310 317 25 143 485
16Res $\mathcal{D}_{1}$ Train 863 775 602 890 43 280 1213
Dev 216 270 237 224 8 66 298
Test 328 342 282 360 25 72 457
$\mathcal{D}_{2}$ Train 857 759 623 1015 50 329 1394
Dev 210 251 221 252 11 76 339
Test 326 338 282 407 29 78 514
Table 5: Statistic information of our two experiment datasets: “#S”, “#T”, “#A”, and “#O” denote the numbers of “Sentences”, “Triplets”, “Aspects”, and “Opinions”; “#S1”, “#S2”, #S3” denote the numbers of sentiments “Positive”, “Neutral” and “Negative”, respectively.
Figure 5: A plot of the hidden word representation, where the dimension is reduced to 2 for convenience of display. “Pretrained” means the model with official released version of weights. “Finetuned” means the model is a finetuned version on ASTE task for certain epochs. “w/o. cl” means the model is trained without contrastive learning loss. “+ cl” means the model is trained with contrastive learning. All the plotted results are from experiment carried on $\mathcal{D}_{1}$ 14Res.
Methods 14Res 14Lap 15Res 16Res
AE OE AOPE AE OE AOPE AE OE AOPE AE OE AOPE
CMLA 81.22 83.07 48.95 78.68 77.95 44.10 76.03 74.67 44.60 74.20 72.20 50.00
RINANTE 81.34 83.33 46.29 77.13 75.34 29.70 73.38 75.40 35.40 72.82 70.45 30.70
Li-unified 81.62 85.26 55.34 78.54 77.55 52.56 74.65 74.25 56.85 73.36 73.87 53.75
GTS 83.82 85.04 75.53 79.52 78.61 65.67 78.22 79.31 67.53 75.80 76.38 74.62
Dual-MRC 86.60 86.22 77.68 80.44 79.90 63.37 75.08 77.52 64.97 76.87 77.90 75.71
ContrASTE (Ours) 86.55 87.04 79.60 82.62 83.41 73.23 86.53 83.05 73.87 85.48 87.06 76.29
$\Delta$F1 -0.05 0.82 1.92 2.18 3.51 7.56 8.31 3.74 6.34 8.61 9.16 0.58
Table 6: F1-score performance on other ABSA tasks: AE, OE, and AOPE. The test is implemented on $\mathcal{D}_{1}$. Results of other models are retrieved from Fei et al. (2022).
Methods 14Res 14Lap 15Res 16Res
P R F1 P R F1 P R F1 P R F1
Pipeline
OTE-MTL Zhang et al. (2020) - - 45.05 - - 59.67 - - 48.97 - - 55.83
[…]
RI-NANTE+ Dai and Song (2019) 31.42 39.38 34.95 21.71 18.66 20.07 29.88 30.06 29.97 25.68 22.30 23.87
[…]
${}^{\natural}$ Peng et al. (2020) 58.89 60.41 59.64 48.62 45.52 47.02 51.7 46.04 48.71 59.25 58.09 59.67
Sequence-tagging
Span-BART Yan et al. (2021) - - 72.46 - - 57.59 - - 60.10 - - 69.98
JET Xu et al. (2020) 67.97 60.32 63.92 58.47 43.67 50.00 58.35 51.43 54.67 64.77 61.29 62.98
MRC based
${}^{\dagger}$ Chen et al. (2021a) 71.32 70.09 70.69 65.12 54.41 59.27 63.71 58.63 61.05 67.74 68.56 68.13
COM-MRC Zhai et al. (2022) 76.45 69.67 72.89 64.73 56.09 60.09 68.50 59.74 63.65 72.80 70.85 71.79
Table-filling
$\mathrm{S}^{3}\mathrm{E}^{2}$ Chen et al. (2021b) 69.08 64.55 66.74 59.43 46.23 52.01 61.06 56.44 58.66 71.08 63.13 66.87
GTS Wu et al. (2020a) 70.92 69.49 70.20 57.52 51.92 54.58 59.29 58.07 58.67 68.58 66.60 67.58
EMC-GCN Chen et al. (2022) 71.85 72.12 71.78 61.46 55.56 58.32 59.89 61.05 60.38 65.08 71.66 68.18
BDTF Zhang et al. (2022) 76.71 74.01 75.33 68.30 55.10 60.99 66.95 65.05 65.97 73.43 73.64 73.51
DGCNAP Li et al. (2023) 71.83 68.77 70.26 66.46 54.34 58.74 62.03 57.18 59.49 69.39 72.20 70.77
LLM-based
GPT 3.5 zero-shot 39.21 56.17 46.18 26.21 40.69 31.88 31.21 52.75 39.21 35.28 59.64 44.34
GPT 3.5 few-shots 44.73 58.87 50.84 29.63 37.69 33.18 37.27 56.42 44.89 43.15 60.75 50.46
GPT 4 zero-shot 27.34 37.13 31.49 16.50 24.41 19.69 25.60 39.22 30.98 28.39 43.64 35.79
GPT 4 few-shots 41.48 52.06 46.17 28.79 39.83 33.42 38.04 58.02 45.96 40.89 62.50 49.44
Ours
ContrASTE 75.87 76.12 76.00 67.45 61.01 64.07 66.84 64.08 65.43 69.38 74.40 71.80
Table 7: Experimental results on $\mathcal{D}_{1}$ Wu et al. (2020a). The best results are highlighted in bold, while the second best results are underlined.
Methods Brief Introduction
Pipeline
OTE-MTL Zhang et al. (2020)
It proposes a multi-task learning framework including two parts: aspect and opinion tagging, along with word-level sentiment dependency parsing. This approach simultaneously extracts aspect and opinion terms while parsing sentiment dependencies using a biaffine scorer. Additionally, it employs triplet decoding based on the aforementioned outputs during inference to facilitate triplet extraction.
Li-unified-R+PD Peng et al. (2020)
It proposes an unified tagging scheme, Li-unified-R, to assist target boundary detection. Two stacked LSTMs are employed to complete aspect-based sentiment prediction and the sequence labeling.
CMLA+C-GCN Wang et al. (2017)
It facilitates triplet extraction by modelling the interaction between the aspects and opinions.
Two-satge Peng et al. (2020)
It decomposes triplet extraction to two stages: 1) predicting unified aspect-sentiment and opinion tags; and 2) pairing the two results from stage one.
RI-NANTE+ Dai and Song (2019)
It adopts the same sentiment triplets extracting method as that of CMLA+, but it incorporates a novel LSTM-CRF mechanism and fusion rules to capture word dependencies within sentences.
Sequence-tagging
Span-BART Yan et al. (2021)
It redefines triplet extraction within an end-to-end framework by utilizing a sequence composed of pointer and sentiment class indexes. This is achieved by leveraging the pretrained sequence-to-sequence model BART to address ASTE.
JET Xu et al. (2020)
It extracts triplets jointly by designing a position-aware sequence-tagging scheme to extract the triplets and capturing the rich interactions among the elements.
MRC-based
Dual-MRC Mao et al. (2021)
It proposes a MRC-based solution for ASTE by jointly training two BERT-MRC models with parameters sharing.
BMRC Chen et al. (2021a)
It introduces a bidirectional MRC (BMRC) framework for ASTE, employing three query types: non-restrictive extraction queries, restrictive extraction queries, and sentiment classification queries. The framework synergistically leverages two directions, one for sequential recognition of aspect-opinion-sentiment and the other for sequential recognition of opinion-aspects-sentiment expressions.
Table-filling
GTS Wu et al. (2020a)
It proposes a novel 2D tagging scheme to address ASTE in an end-to-end fashion only with one unified grid tagging task. It also devises an effective inference strategy on GTS that utilizes mutual indication between different opinion factors to achieve more accurate extraction.
Double-encoder Jing et al. (2021)
It proposes a dual-encoder model that capitalizes on encoder sharing while emphasizing differences to enhance effectiveness. One of the encoders, referred to as the pair encoder, specifically concentrates on candidate aspect-opinion pair classification, while the original encoder retains its focus on sequence labeling.
$\mathrm{S}^{3}\mathrm{E}^{2}$ Chen et al. (2021b)
It represents the semantic and syntactic relationships between word pairs, employs GNNs for encoding, and applies a more efficient inference strategy.
EMC-GCN Chen et al. (2022)
It employs a biaffine attention module to embed ten types of relations within sentences, transforming the sentence into a multi-channel graph while incorporating various linguistic features to enhance performance. Additionally, the method introduces an effective strategy for refining word-pair representations, aiding in the determination of whether word pairs are a match or not.
LLM-based
zero-shot
Performing aspect-based sentiment analysis using an LLM. The specific method involves inputting a prompted sentence and directly outputting the corresponding [A, O, S] triplets. An example of the text given to the LLM, with the prompt added, is as follows: "Perform aspect-based sentiment analysis on the provided text and return triplets as [Aspect, Opinion, Sentiment]. You only need to provide the triplets, no additional explanations are required. The provided text: {sentence}"
few-shots
Building upon the zero-shot method, a small number of examples from the training set are added to the prompted sentence: "Perform aspect-based sentiment analysis on the provided text and return triplets as [Aspect, Opinion, Sentiment]. For example: input: {train sentence} output: {train triplets}, … (some other examples). You only need to provide the triplets, no additional explanations are required. The provided text: {sentence}"
Table 8: Baselines methods with brief introduction.
Sentence Ground Truth Predictions
GPT-3.5 zero-shot
GPT-3.5 few-shots
GPT-4 zero-shot
GPT-4 few-shots
It is a cozy place to go with a couple of friends.
[place, cozy, positive]
[place, cozy, positive],
[friends, couple of, positive]
[place, cozy, positive]
[place, cozy, positive]
[going with friends, a couple of, positive]
[place, cozy, positive]
This place has the best sushi in the city.
[sushi, best, positive]
[sushi, best, positive],
[place, best, positive], [city, best, positive]
[sushi, best, positive]
[sushi, the best, positive],
[place, This, neutral],
[city, in the, neutral]
[sushi, best, positive]
Disappointingly, their wonderful Saketini has been taken off the bar menu.
[Saketini, wonderful, positive],
[bar menu, Disappointingly, negative]
[Saketini, wonderful, positive],
[bar menu, taken off, negative]
[Saketini, taken off, negative]
[Saketini, has been taken off the bar menu, negative],
[bar menu, wonderful, positive]
[Saketini, wonderful, positive],
[bar menu, taken off, negative]
I went in one day asking for a table for a group and was greeted by a very rude hostess.
[hostess, rude, negative]
[hostess, rude, negative]
[hostess, rude, negative]
[table for a group, asking, neutral],
[hostess, very rude, negative]
[table for a group, asking, neutral],
[hostess, very rude, negative]
But make sure you have enough room on your credit card as the bill will leave a big dent in your wallet.
[bill, big, negative]
[credit card, enough room, positive],
[bill, big dent, negative],
[wallet, big dent, negative]
[credit card, enough room, negative],
[bill, big dent, negative],
[wallet, big dent, negative]
[room on your credit card, enough, positive],
[bill, will leave a big dent in your wallet, negative]
[bill, big dent, negative]
Table 9: In summary, there are several challenges observed in the performance of GPT models concerning triplets. Firstly, there is a prominent issue of "hard" matching, where GPT models tend to introduce additional modifiers or adverbs in the opinion component, leading to a lack of exact correspondence. Secondly, during zero-shot inference, GPT models tend to generate multiple predicted triplets, resulting in decreased precision. This behavior particularly hampers the precision of the model’s predictions. Thirdly, inconsistencies arise in handling triplets involving structures such as [A, O1 and O2, S] and [A, O1, S], [A, O2, S]. This inconsistency is challenging to mitigate due to its dependence on annotation practices and conventions. Upon closer examination, the issues observed do not appear to be as pronounced as indicated by the evaluation metrics. Rather, they often manifest as cases where the general idea is correctly captured, but the precise format or phrasing does not align perfectly. Notably, the performance of GPT-4 deteriorates due to its occasional tendency to not merely "extract" fragments from sentences but to generate its own summarizations. Consequently, evaluating against triplets that originate solely from annotated sentences poses a challenge in achieving alignment. Furthermore, GPT-4 exhibits a proclivity for extracting longer sequences of words as aspects or opinions, while GPT-3.5 tends to produce shorter sequences that better conform to typical annotation scenarios.